OpenStudy (anonymous):

summation (a sub k+1.x^k) + summation (a sub k. x^k+1) = a sub 1 + summation ( a sub k+1 + a sub k-1 ). x^k....How to verify the given equation urgent help needed :(

OpenStudy (anonymous):

can u write it using mathematical symbols?

OpenStudy (anonymous):

\[\sum_{k=0}^{\infty} a _{k+1}x ^{k} + \sum_{k=0}^{\infty}a _{k}x ^{k+1} = a _{1}+\sum_{k=1}^{\infty} (a _{k+1}+a _{k-1})x ^{k}\]

OpenStudy (anonymous):

Here it goes. Please help me :(